OpenStudy (anonymous):

Find the three formulas for cos 2(theta).

hartnn (hartnn):

i have 4

OpenStudy (anonymous):

you do?

hartnn (hartnn):

but i''l give 3 only you know any 1?

OpenStudy (anonymous):

Cos2 Theta= Cos(Theta +theta).

hartnn (hartnn):

no..... you know this one cos 2x = cos^2x-sin^2x

OpenStudy (anonymous):

yes

hartnn (hartnn):

now use sin^2x+cos^2x=1

hartnn (hartnn):

so whats cos^2x from there?

OpenStudy (anonymous):

1-sin squared theta?

hartnn (hartnn):

yes, plug that in cos 2x formula

OpenStudy (anonymous):

1-sin squared theta- sin squared theta?

OpenStudy (anonymous):

1-2sin squared theta?

hartnn (hartnn):

yes., thats another formula :)

hartnn (hartnn):

now last

hartnn (hartnn):

whats cos^2x from there ......from sin^2x +cos^2x=1

OpenStudy (anonymous):

cos squared theta=1-sin squared theta?

OpenStudy (anonymous):

or ehh 1- sin squared theta=1-sin squared theta

hartnn (hartnn):

thats done right?

OpenStudy (anonymous):

1=1

hartnn (hartnn):

sin^2 x =1-cos^2x put this in cos 2x formula

OpenStudy (anonymous):

hmm...

OpenStudy (anonymous):

cos squared theta+sin squared theta=1 but isn't the formula...

OpenStudy (anonymous):

but could you do sin squared theta=1-1-sin squared theta?

OpenStudy (anonymous):

2 sin squared theta=1-1

OpenStudy (anonymous):

that would equal zero though

hartnn (hartnn):

i mean sin^2 x =1-cos^2x in cos^2x-sin^2x

hartnn (hartnn):

cos^2x - (1-cos^2x) =...?

OpenStudy (anonymous):

2 cos squared theta-1?

hartnn (hartnn):

there you go, 3rd formula

OpenStudy (anonymous):

ty, son goku

hartnn (hartnn):

welcome ^_^